OpenStudy (anonymous):

what are true about the graph of y > x^2 + 4x + 4? 1. The parabola opens up. 2. The line of symmetry is y = -2. (or both?)

OpenStudy (anonymous):

I think it is y= not y>

OpenStudy (anonymous):

no, thats the right question..

OpenStudy (anonymous):

I dont think it is a parabola then

OpenStudy (anonymous):

so theyre both not true?

OpenStudy (anonymous):

It's a parabola. You just have an inequality. The inequality is irrelevant when it comes to just describing the shape. It's relevant in the solution set.

OpenStudy (anonymous):

The dominant term (the quadratic or ^2 term) is not negative, so it will open up.

OpenStudy (waleed_imtiaz):

The first option is true that the parabola will open upwards...... and I am not confirm with the second.

OpenStudy (anonymous):

The line of symmetry will be about the vertex given by: \[-b/2a\]

OpenStudy (anonymous):

Just be careful about how they're wording it, y = -2 or x = -2 ;)

OpenStudy (anonymous):

hmm so the second one is true?

OpenStudy (anonymous):

Here b = 4 (the linear term), a = 1 (the quadratic term) by -b/2a we have -4/(2 * 1) = -2